OpenStudy (anonymous):

Solve the equation. 3(x – 9) = 30 x = 1 x = 13 x = 19 x = 57

OpenStudy (anonymous):

Substitute the values into the equation

OpenStudy (anonymous):

You are not supposed to give them answer. You have to help so they understand

OpenStudy (monkey202):

best way to get the answer is to substitute them in

OpenStudy (anonymous):

it would be 3(x-9)=30 3x-27=30 +27=+27 30+27 = 57 3x=57

OpenStudy (anonymous):

you have to solve the rest of the problem